A parallelogram has vertices L(-2,5), M(3,3), N(1,0). What are possible coordinates for its fourth vertex? EXPLAIN
Answer by Fombitz(32388)


Look at the relationship between points (1,0) and (3,3).
Delta X=3-1=2
Delta Y=3-0=3
Point (-2,5) must maintain the same relationship with the other point.
Two points can do this,
(-2+2,5+3)=(0,8)
(-2-2,5-3)=(-4,2)